OpenStudy (anonymous):

Simplify the difference b^2-2b-8/b^2+b-2 subtract 6/b-1

OpenStudy (anonymous):

factorize \[(b ^{2}-2b-8)/(b ^{2}+b-2)\] \[(b ^{2}-2b-8)/(b ^{2}+b-2) = (b=4)(b+2)/(b+2)(b-1)\]

OpenStudy (anonymous):

from there, you can do the subtraction

OpenStudy (anonymous):

it's supposed to be (b-4) and not (b=4) sorry. typo mistake

OpenStudy (anonymous):

so what will my answer be I can do it from there but I wanna make sure I get the right answer! Thank u for ur help I appreciate it!

OpenStudy (anonymous):

you will get (b-10)/(b-1) (:

OpenStudy (anonymous):

Thank u very very much!!!